Quadratic Equation
2y^2 - 3y - 6 = 0
Textbook Elementary and Intermediate Algebra. Mark Dugopolski
ISBN: 13978-0-07-326862-0

Answer by mukhopadhyay(490)

2y^2 - 3y - 6 = 0... (multi 1/2)
=>y^2-(3/2)y=3 ... [add (3/4)^2]
=>(y-3/4)^2 = 3+9/16
=>(y-3/4)^2 = 57/16
=>(y-3/4)^2 = [sqrt(57)/4]^2
=>y=(3/4)+[sqrt(57)/4] or y=(3/4)-[sqrt(57)/4]